### v1

# Strategizing with AI: Insights from a Beauty Contest Experiment

Abstract
A Keynesian beauty contest is a wide class of games of guessing the most popular strategy among other players. In particular, guessing a fraction of a mean of numbers chosen by all players is a classic behavioral experiment designed to test iterative reasoning patterns among various groups of people. The previous literature reveals that the level of sophistication of the opponents is an important factor affecting the outcome of the game. Smarter decision makers choose strategies that are closer to theoretical Nash equilibrium and demonstrate faster convergence to equilibrium in iterated contests with information revelation. We replicate a series of classic experiments by running virtual experiments with modern large language models (LLMs) who play against various groups of virtual players. We test how advanced the LLMs’ behavior is compared to the behavior of human players. We show that LLMs typically take into account the opponents’ level of sophistication and adapt by changing the strategy. In various settings, most LLMs (with the exception of Llama) are more sophisticated and play lower numbers compared to human players. Our results suggest that LLMs (except Llama) are rather successful in identifying the underlying strategic environment and adopting the strategies to the changing set of parameters of the game in the same way that human players do. All LLMs still fail to play dominant strategies in a two-player game. Our results contribute to the discussion on the accuracy of modeling human economic agents by artificial intelligence.
Keywords: beauty contest, guess the number, dominant strategy, AI, LLM
JEL codes: C99, D90, C72
1 Introduction
The recent appearance of large language models (LLMs) has resulted in numerous attempts to substitute humans with generative agents in various settings (see, for example, Park et al. (2023)). The motivation is simple: in many economic activities, the use of algorithms sooner or later will be cheaper and more productive than the use of labor. Still, at this point, it is not clear to what extent LLM can simulate the human’s behavior. Several papers investigated the differences between LLM decisions and the participants of economic experiments. In the recent manuscript, Horton (2023) replicates several classic experiments with LLM players and advocates the use of LLMs as models for ordinary economic agents. Our paper contributes to this strand of literature by studying the behavior of LLMs in the classic Guess the number game. This game belongs to a wider class of Keynesian beauty contest experiments and is of particular importance because it is designed to study level-$k$ reasoning, the ability of a player to make a sequence of conclusions as a function of their level of sophistication. To what extent LLM behaves like a human in making sequential conclusions? To answer this question, we replicate a series of well-known experiments with human participants by asking LLMs to play against the same groups of competitors.
The rules of the Guess the number game are as follows. A group of $n$ players simultaneously and independently choose a number between 00 and $100$. Denote by $m$ the mean of all strategies played. A player whose number is the closest to $pm$, where $p>0$ is the predetermined constant known to all players before the game, wins. In case of a tie, all tied players get the corresponding share of the prize. For $p\in(0,1)$, there is a unique Nash equilibrium in the model where all players choose 00. In a particular case of $n=2$, choosing 00 is a weakly dominant strategy.
Multiple experiments show that in the Guess the number game people, in general, do not play Nash equilibrium. In the pioneering experimental paper, Nagel (1995) demonstrated that in sessions with $p=1/2$ and $p=2/3$ no subject chose 00 and only 6 percent chose numbers below $10$. However, in the iterated game the strategies converged to Nash equilibrium from period to period, after the participants learned statistics from the previous rounds (Nagel, 1995). If one uses the median of the chosen numbers instead of the mean, results do not change much in a one-shot game but in the iterated game convergence to 00 is faster in the median variant compared to the mean variant (Duffy & Nagel, 1997). Switching to the maximum instead of the mean or the median increases the chosen strategies (Duffy & Nagel, 1997).
In a particular case of $n=2$, one could possibly anticipate a significant share of players choosing 00, a weakly dominant strategy. However, this is not the case. Only $10\%$ of undergraduate students and $37\%$ of the audience of economics or psychology decision-making conferences chose 00 (Grosskopf & Nagel, 2008). Also, the mean of the numbers chosen by the professionals (22) is lower than the mean of the numbers chosen by the students (36). A higher number of participants $n=18$ leads to a lower mean both for professionals (19) and students (29). However, in the case of professionals, this difference is not statistically significant (Grosskopf & Nagel, 2008).
Several theoretical models explaining the non-equilibrium behavior were proposed in the literature. Most of these models deal with the notion of bounded rationality when players are rational only to some extent; the degree of rationality is associated with the sophistication of a player. A dynamical model where the players choose one of the step-$k$ behavioral rules, learn the results of the experiment, and choose more successful rules in the next iterations, was presented and estimated in Stahl (1996). A further extension of the set of possible behavioral strategies is discussed in Stahl (1998). Ho et al. (1998) builds the bounded rationality models based on the iterative deletion of dominated strategies and iterated best replies to previously played actions. It appears that many participants of experiments are using iterated best response arguments. Namely, Bosch et al. (2002) describes an experiment organized by The Financial Times where 64% of players indeed explained that they exploited the best responses to the revealed statistics. Note that playing iterative best response to the previous iteration of the game does not lead to the best response to other players’ strategies in the new iteration (Breitmoser, 2012). Weber (2003) demonstrated that the feedback from organizers plays a key role in the speed of convergence to Nash equilibrium: in the absence of the feedback, the numbers also decreased but at a lower rate. Advice from a peer participant has an even stronger effect on the performance than pure statistics provided by the organizers (Kocher et al., 2014). The authors of the latter paper relate it to the limited ability of players to analyze statistical data.
Cognitive ability is also an important determinant of the outcome of the Guess the number game. Brocas & Carrillo (2020) designed a variant of the Guess the number game and demonstrated that the equilibrium behavior increases significantly between 5 and 10 years of age (from 17.9% to 61.4%) and stabilizes afterwards. Back to the claasic variant of the game, players with higher scores in a cognitive ability test choose lower numbers (Burnham et al., 2009). Mixed evidence was reported in Brañas-Garza et al. (2012): the better performance in the CRT test that measures cognitive reflectiveness is associated with lower numbers in the Guess the number game, whereas the outcome of the Raven test measuring visual reasoning and analytic intelligence surprisingly was not associated with the successful performance in the Guess the number game.
Another evidence that the level of players’ sophistication matters, comes from experiments where teams consisting of several players played instead of single players. The strategies of teams of 2 players do not differ significantly from the strategies chosen by individual players (Sutter, 2005). At the same time, teams consisting of 3 and 4 players outperform individual players (Kocher & Sutter, 2005, Sutter, 2005).
One could hypothesize that emotions affect the players’ decisions by diminishing the ability to perform deep analysis of the game. However, the evidence differs for various conditions. Players experiencing stress during the game indeed choose higher numbers (Leder et al., 2013). Angry participants of the experiment have a lower level of reasoning compared to the control group (Castagnetti et al., 2023). At the same time, sadness has little effect on the players’ strategies (Castagnetti et al., 2023).
It appears that framing of the problem also matters. Hanaki et al. (2019) considered two variants of the Guess the number game. In the first one, the players’ responses are strategic complements, while in the second one their responses are strategic substitutes. Theoretical Nash equilibrium is the same in both variants. However, the authors demonstrate that the strategic environment effect manifests itself for sufficiently large or uncertain groups of players: the players’ responses begin to diverge starting from 5-players games.
For more experimental and theoretical results on Guess the number games, we refer the reader to one of the surveys (Nagel, 2008, Nagel et al., 2017).
Testing LLMs with the Guess the number game offers a dual benefit. By comparing model outputs with well-established human data, we can assess whether LLMs capture the bounded rationality and iterative reasoning typical of human decision-making. At the same time, any systematic differences highlight limitations in LLM strategic reasoning, providing valuable insights for refining these models and for their deployment in economic applications.
By performing a series of experiments, we aimed to investigate the decision-making process and strategy formulation of the modern LLMs playing against different groups of virtual players. We want to identify to what extent LLMs behave like a human player and whether LLMs can successfully identify the other players’ level of sophistication. To move forward on this path, we break our task down into several intermediate questions.
Q1. Does LLM recognize the rules of the game and act in accordance with the rules?
This is the simplest but necessary test of LLM abilities. A negative answer would mean that further investigations are pointless.
Q2. Does LLM recognize the strategic context of the game?
It is important to understand whether LLM takes into account the strategies of other players when choosing their own strategy. The negative answer would make us think of LLM as a very simple, unsophisticated player.
Q3. Are LLM’s decisions in line with the expected comparative statics with respect to the parameters of the experiment?
Theoretical models of bounded rationality and the empirical evidence predict that more sophisticated players behave closer to Nash equilibrium in the Guess the number game. We aim to test how LLM responds to changes in parameters of the game. Failing to fulfill the expectations will be perceived as a disappointing signal about the LLM abilities.
Q4. Can LLM find an analytical solution to the game?
This is an interesting question per se because neither positive nor negative answer would make the comparison of human and LLM strategies less meaningful. Recall that some of the previous experiments included those participants who are definitely unfamiliar with dominant strategies and Nash equilibrium concepts, whereas other experiments included game theorists who are presumably able to find the theoretical equilibrium in the Guess the number game.
Q5. Can LLM correctly use the analytical solution when choosing their strategy?
In case of a positive answer, for the game of $n=2$ players we would expect that LLM always plays a weakly dominant strategy. For the $n>2$ case, we do not infer too strong conclusions because dealing with the real-world competitors requires the correct evaluation of their complexity rather than following the theoretical predictions.
Q6. Do LLMs perform differently?
On the one hand, treating LLMs as black boxes means that we expect potentially different outcomes depending on what is inside those boxes. On the other hand, some tasks can be so simple (for example, asking to find the sum 2+2) that different models would provide the same output.
Q7. Are LLM’s strategies similar to strategies played by human players?
Though this question is the most important for us, the nuances that may arise in questions Q1Q6 could affect the interpretation of the results. Therefore, we do not restrict our attention to Q7 solely.
We experimented with the main LLMs that were available in 2024: GPT-4o, GPT-4o-Mini, Gemini-1.5-flash, Claude-3.5-Sonnet (20240620), Llama-3.1-8B-Instruct-Turbo. Apart from the comparison of human versus LLMs’ behavior, we discuss the results of LLMs’ behavior differences.
We found that LLMs tend to see competitors as more sophisticated agents than human players do. Note that this is not a general property of LLMs. For example, in a money request game LLMs are found to be less sophisticated than human players (Gao et al., 2024).
After the initial publication of our working paper (Dagaev et al., 2024), we found several new relevant working papers that we would like to mention. Lu (2024) also runs beauty contest experiments with AI players. Whereas the latter paper focuses on the iterated standard beauty contest game and dynamics of the LLMs decisions, our paper replicates a wide class of one-shot experiments with different groups of competitors. Guo et al. (2024) introduced EconArena environment that allows to organize various competitions between AI models. A beauty contest game is one of the first games that was implemented on EconArena.
The rest of the paper is organized as follows. In Section 2, we describe the methodology of our research. Section 3 presents main results that are further discussed in Section 4. Section 5 concludes.
2 The Experimental Design
We run the experiments with the following LLMs:
  1. 1.
    GPT-4o — an OpenAI model released in May 2024. OpenAI describes it as a "step towards much more natural human-computer interaction" [^1] OpenAI official website: https://openai.com/index/hello-gpt-4o/. Retrieved December 22, 2024 because it accepts multimodal information and quickly analyzes it to give a response.
  2. 2.
    GPT-4o mini — a smaller version of GPT-4o, described by OpenAI as the "most cost-efficient small model". [^2] OpenAI official website: https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/. Retrieved December 22, 2024. OpenAI also claims that it outperforms other small models on such tasks as reasoning, math and coding, multimodal reasoning.
  3. 3.
    Gemini-1.5-flash — a model from Deepmind, described as lightweight and optimized for tasks requiring speed and efficiency. [^3] Deepmind official website: https://deepmind.google/technologies/gemini/flash/. Retrieved December 22, 2024.
  4. 4.
    Claude-3.5-Sonnet (20240620) — an Anthropic-made model. Creators claim that it "sets new industry benchmarks for graduate-level reasoning (GPQA), undergraduate-level knowledge (MMLU), and coding proficiency (HumanEval)", making it a good choice for our experiments. [^4] Anthropic official website: https://www.anthropic.com/news/claude-3-5-sonnet. Retrieved December 22, 2024.
  5. 5.
    Llama-3.1-8B-Instruct-Turbo — the smallest of the Meta Llama 3.1 model 2024 releases. As Meta claims, it outperforms such models as Gemma 2 and Mistral 7B Instruct in a range of tasks (e.g., coding, reasoning, math). [^5] Meta official website: https://ai.meta.com/blog/meta-llama-3-1/. Retrieved December 22, 2024.
There are several reasons for such a choice of LLMs. First, we want to know whether our results are robust to differently designed and sized LLMs: our selected models are taken from several companies (Google Deepmind, Meta, OpenAI, Anthropic) and possess different approaches and architectures. Second, we experiment with both open source (Llama) and proprietary models (GPT, Claude, Gemini) that further enhances robustness of our results. Third, selected models share top positions in various benchmarks, allowing us to deal with the state-of-the-art LLMs. For example, they appear at high positions on the Chatbot Arena LLM Leaderboard. [^6] https://lmarena.ai/?leaderboard. Retrieved December 22, 2024.
The experiments were structured into 16 distinct scenarios, each characterized by a combination of factors including the type of aggregate statistic used to determine the winning number (Function), the target percentage ($p$) of the aggregate statistic, the number of players involved ($n$), and the composition of the opponent group (Opponents). These factors are identical to the settings of classic experiments with real people reported in previous literature (see Table 1 for the summary).
Scenario Original paper n p Function Opponents
1 Nagel (1995) 18 1/2 Mean Undergraduate students of various faculties
2 18 2/3 Mean Undergraduate students of various faculties
3 Duffy & Nagel (1997) 16 1/2 Mean Undergraduate students
4 16 1/2 Median Undergraduate students
5 16 1/2 Maximum Undergraduate students
6 Grosskopf & Nagel (2008) 2 2/3 Mean First year undergraduate students majoring in economics
7 2 2/3 Mean Audience of economics or psychology-decision making conferences
8 18 2/3 Mean First year undergraduate students majoring in economics
9 18 2/3 Mean Audience of game theory conferences
10 Brañas-Garza et al. (2012) 24 1/2 Mean individuals with high CRT score
11 24 1/2 Mean individuals with low CRT score
12 24 2/3 Mean individuals with high CRT score
13 24 2/3 Mean individuals with low CRT score
14 Castagnetti et al. (2023) 3 0.7 Mean individuals experiencing anger
15 3 0.7 Mean individuals experiencing sad emotions
16 3 0.7 Mean individuals experiencing neither anger nor sad emotions
Table 1: Summary of experiments replicated in this paper with an AI player. CRT = cognitive reflection test.
In each scenario, we simulate one of the human participants using the LLM. Our goal is to explore the effects of specified conditions on the model’s responses and compare the AI’s strategies with human strategies, given all else equal. The procedure commences by addressing the application programming interface (API) with prompts containing information regarding the experimental conditions including a description of the opponents, the values of $p$ and $n$, and the specific aggregate function. Following the delivery of prompts, the responses generated by LLM were systematically recorded, capturing not only the model’s guess regarding the number but also its underlying reasoning and conclusions. To account for variability in the responses, this procedure was iterated 50 times for each scenario. Each query was treated as a new instance to avoid any learning effects. Additionally, zero-shot prompting was employed, with no examples provided to the model, in order to capture its raw reasoning while minimizing potential biases from prior interactions or fine-tuning on related tasks. This approach was chosen to closely replicate the experimental conditions with human players described above.
The following prompt for LLMs was formulated to simulate Guess the number game taking into account the parameters of the game described above:
‘You are playing the game. In this game, $n$ players including you simultaneously choose an integer between 0 and 100, inclusive. The winner of the game is the player whose number is closest to $p$ of the $function$ of all of the numbers chosen. You are playing against the following opponents: $Opponents$. You are in the same group as the $Opponents$, you possess the same skills, knowledge, level of intelligence, education, etc. What number will you choose, and why? Provide your answer in a JSON format with two keys: "answer_text" (which contains the full text of the answer including the reasoning and concluded by your unambiguously chosen number) and "answer_number" (which contains only your unambiguously chosen number).’
Below, we provide an example of the GPT-4o-Mini text response to the prompt that was collected into a "answer_text" field in scenario 6 ($n=2$, $p=2/3$, mean is the aggregate function, and the opponents are first year undergraduate students majoring in either economics, political science, law, medicine or humanities with no formal training in game theory).
‘In this game, the goal is to choose a number that is closest to 2/3 of the average of all numbers chosen by the players. If all players in the game are first-year undergraduate students majoring in various fields with no formal training in game theory, they may not fully grasp the strategic implications of this game. Many players might tend to choose numbers that are higher due to lack of understanding of the average, or they may attempt to guess what others will choose without coordinated strategizing. Since the maximum number is 100, the average of all choices might lean towards the higher end. However, to achieve the objective of getting closest to 2/3 of that average, I should anticipate that the average could be around the middle range, say 50, which would make 2/3 of that average about 33.33. Given that my opponents are not adept at game theory, they may tend to gravitate towards higher numbers. Thus, I will choose a number lower than 33.33 to ensure I am closer to this target. After considering these dynamics, I will choose the number 25, as it positions me well to be closest to 2/3 of a potentially inflated average.’
In the same experiment, the model answered with a ‘25’ in a "answer_number" field.
All answers in all scenarios respected the format of data provision requested in the prompt. We accepted all answers of the models irrespectively of the potential logical drawbacks in explanations. Therefore, we have 50 answers from each of 5 LLMs in each of 16 scenarios.
3 Results
This section presents the results of our analysis, comparing the performance of various AI models across 16 experimental settings derived from five previous studies of beauty contest games. We focus on the LLMs’ alignment with theoretical predictions and their consistency with observed human behavior by comparing model-generated decisions to those of human participants in the original experiments. Additionally, we assess how the models’ strategies diverge or converge by comparing them to one another. Our findings are organized in five tables, each corresponding to a different original study, and together they illustrate the extent to which AI-driven agents replicate human-like reasoning and adapt to strategic environments.
Scenario $n$ $p$ Function Opponents Paper Model Model Model $MM-PM$ $t$-stat $p-$value
mean mean st. dev.
($PM$) ($MM$)
18 1/2 mean undergraduate students 27.05 Gemini Flash 13.42 19.47 -13.63 -4.95 0.000
Claude Sonnet 10.28 3.12 -16.77 -37.96 0.000
1 GPT-4o mini 14.72 9.34 -12.33 -9.34 0.000
GPT-4o 16.94 4.63 -10.11 -15.45 0.000
Llama 42.14 11.26 15.09 9.47 0.000
18 2/3 mean undergraduate students 36.73 Gemini Flash 9.82 13.84 -26.91 -13.75 0.000
Claude Sonnet 17.76 6.45 -18.97 -20.78 0.000
2 GPT-4o mini 17.98 8.51 -18.75 -15.58 0.000
GPT-4o 21.16 4.67 -15.57 -23.57 0.000
Llama 35.68 11.17 -1.05 -0.66 0.509
Table 2: Replication results for Nagel (1995) with a LLM player. In Nagel (1995), $n$ varies from 15 to 18 in different sessions. In our experiments, we fixed the number of players at 18 and assumed that the marginal effect of one additional player in the group of 15–18 players is low.
Table 2 shows replicated results for the pioneering Nagel (1995) experiment, which explores strategic reasoning among undergraduate students from various faculties. The table reports results for two scenarios: one with a target fraction $p=1/2$ and another with $p=2/3$, both applied to the average of all responses. In both scenarios, the AI agents demonstrate a tendency to produce guesses closer to zero, the Nash equilibrium strategy, compared to the human participants’ averages reported in the paper. For instance, in the $p=1/2$ case, the mean guesses of all AI models except Llama are significantly lower than the human mean of 27.05, with differences ranging from -10.11 to -16.77, all statistically significant at any reasonable level. For $p=2/3$, all AI models tend to play closer to zero, with model means ranging from 9.82 (Gemini Flash) to 35.68 (Llama), compared to the human mean of 36.73. Interestingly, the Llama model consistently deviates from other models, producing higher guesses that are closer to human participants in the $p=2/3$ case and overshooting significantly in the $p=1/2$ case. These variations are probably due to differences in how the models interpret strategic reasoning under bounded rationality and reflect the diversity of decision-making paradigms across LLMs.
Table 3 replicates findings from the Duffy & Nagel (1997) experiment which compares decision-making strategies using three different aggregation methods: the mean, the median, and the maximum. The participants in the original study were undergraduate students. The paper reported no significant differences between the strategies in the mean and the median games while in the maximum game people choose significantly higher numbers than in either the mean or median games. For the median aggregation function, the AI agents display a range of mean guesses, with Claude Sonnet providing lower mean guess (14.08), closer to theoretical expectations compared to Llama, which shows the highest mean (32.00). On the other hand, for the mean aggregation function, the results reveal more variation, with Gemini Flash and Claude Sonnet yielding significantly lower guesses (9.08 and 11.00, respectively), compared to the higher mean of 43.30 observed for Llama. While the paper’s original findings reported no significant differences between the median and mean aggregation methods, the AI agents exhibit some divergence in their performance, reflecting differences in how the models process these aggregation rules. Once again, Llama is a very different agent compared to other LLMs.
Scenario $n$ $p$ Function Opponents Paper result Model Model mean Model st. dev
16 1/2 mean undergraduate students Difference with median game is not significant Gemini Flash 9.08 15.62
Claude Sonnet 11.00 2.60
3 GPT-4o mini 16.60 7.64
GPT-4o 15.72 5.06
Llama 43.30 12.35
16 1/2 median undergraduate students Difference with mean game is not significant Gemini Flash 26.89 18.64
Claude Sonnet 14.08 5.62
4 GPT-4o mini 21.86 5.50
GPT-4o 23.44 3.56
Llama 32.00 15.53
16 1/2 maximum undergraduate students Significantly higher numbers Gemini Flash 31.9 28.89
Claude Sonnet 32.54 9.19
5 GPT-4o mini 24.24 6.36
GPT-4o 30.2 6.24
Llama 46.48 8.49
Table 3: Replication results for Duffy & Nagel (1997) with a LLM player.
Table 4 presents our replication of the experiment reported in Grosskopf & Nagel (2008). In the original study, multiple audiences were involved, allowing for an examination of how domain knowledge and familiarity with strategic thinking might influence guesses in a guesses in a beauty contest.
In Scenarios 6 and 7 (both two-player games), the original human means were 35.57 for undergraduate students and 21.73 for conference audiences (Economics/Psychology-decision making). The first four AI models (Gemini Flash, Claude Sonnet, GPT-4o Mini, GPT-4o) all guess significantly lower than these human means. In contrast, Llama’s guesses exceed human averages by approximately 9 to 24 points. A similar pattern appears in Scenarios 8 and 9, where 18-player games yielded human means of 29.31 (undergraduates) and 18.98 (game theory conference audiences). Here again, Gemini Flash, Claude Sonnet, GPT-4o Mini, and GPT-4o consistently guess well below human averages. Llama again overshoots human averages by 12.51 to 18.78. All of these differences are statistically significant at conventional levels. Overall, the results of replication of this paper align with our earlier observations that most AI models exhibit guesses closer to the Nash equilibrium (zero), even when human participants themselves may be relatively sophisticated (e.g., game theory conference attendees).
Scenario $n$ $p$ Function Opponents Paper Model Model Model $MM-PM$ $t-$stat $p-$value
mean mean st. dev.
($PM$) ($MM$) st. dev.
2 2/3 mean First year undergraduate students majoring in Economics 35.57 Gemini Flash 25.52 22.30 -10.05 -3.19 0.000
Claude Sonnet 25.48 7.55 -10.09 -9.45 0.000
6 GPT-4o mini 21.34 5.04 -14.23 -19.96 0.000
GPT-4o 24.14 4.20 -11.43 -19.25 0.000
Llama 44.70 13.50 9.13 4.78 0.008
2 2/3 mean Audience of economics or psychology-decision making conferences 21.73 Gemini Flash 10.36 16.93 -11.37 -4.75 0.000
Claude Sonnet 14.20 4.02 -7.53 -13.24 0.000
7 GPT-4o mini 10.30 10.42 -11.43 -7.76 0.000
GPT-4o 15.40 9.99 -6.33 -4.48 0.014
Llama 45.52 12.74 23.79 13.21 0.000
18 2/3 mean First year undergraduate students majoring in Economics 29.31 Gemini Flash 13.50 16.26 -15.81 -6.87 0.000
Claude Sonnet 22.78 6.13 -6.53 -7.53 0.000
8 GPT-4o mini 22.64 6.59 -6.67 -7.16 0.000
GPT-4o 23.70 4.12 -5.61 -9.64 0.000
Llama 41.82 10.72 12.51 8.25 0.000
18 2/3 mean Audience of game theory conferences 18.98 Gemini Flash 2.74 11.22 -16.24 -10.24 0.000
Claude Sonnet 7.40 5.51 -11.58 -14.85 0.000
9 GPT-4o mini 4.10 8.02 -14.88 -13.12 0.000
GPT-4o 5.90 8.28 -13.08 -11.17 0.000
Llama 37.76 12.01 18.78 11.06 0.000
Table 4: Replication results for Grosskopf & Nagel (2008) with a LLM player.
Replication results for Brañas-Garza et al. (2012) experiment are reported in Table 5. The latter paper examines the influence of cognitive reflection on decision-making in strategic games. The original study reported that individuals with higher Cognitive Reflection Test (CRT) scores tend to choose lower numbers compared to those with lower CRT scores.
Across both values of $p$, Gemini Flash, Claude Sonnet, GPT-4o Mini, and GPT-4o display lower guesses under the high-CRT scenarios (Scenarios 10 and 12) compared to the low-CRT ones (Scenarios 11 and 13), aligning well with the human data from Brañas-Garza et al. (2012). By contrast, Llama’s guesses remain relatively high in all conditions, echoing its tendency to overshoot in previous experiments. Nonetheless, even Llama shows a small but consistent upward shift from the high- to the low-CRT scenarios, broadly reflecting the original study’s observed relationship between CRT scores and numerical guesses.
Model $n$ $p$ Function Opponents Paper result Model Model mean Model st. dev.
24 1/2 mean individuals with high cognitive reflection test (CRT) score Lower numbers compared to players with low CRT scores Gemini Flash 4.90 16.88
Claude Sonnet 2.96 2.35
10 GPT-4o mini 7.62 9.18
GPT-4o 9.82 6.60
Llama 47.94 7.07
24 1/2 mean individuals with low cognitive reflection test (CRT) score Higher numbers compared to players with high CRT scores Gemini Flash 21.46 15.92
Claude Sonnet 26.94 10.41
11 GPT-4o mini 19.9 5.70
GPT-4o 22.46 3.68
Llama 46.86 10.27
24 2/3 mean individuals with high cognitive reflection test (CRT) score Lower numbers compared to players with low CRT scores Gemini Flash 2.54 7.93
Claude Sonnet 6.74 5.07
12 GPT-4o mini 14.5 9.73
GPT-4o 9.54 8.89
Llama 38.46 9.89
24 2/3 mean individuals with low cognitive reflection test (CRT) score Higher numbers compared to players with high CRT scores Gemini Flash 18.86 18.88
Claude Sonnet 41.94 11.92
13 GPT-4o mini 24.2 5.30
GPT-4o 28.94 4.78
Llama 41.06 12.30
Table 5: Replication results for Brañas-Garza et al. (2012) with a LLM player.
Table 6 replicates the experiment from Castagnetti et al. (2023) who investigate the impact of emotions, particularly anger and sadness, on decision-making in strategic games. The original paper concluded that individuals experiencing anger performed less optimally compared to a control group. At the same time, players who experience sad emotions, do not play significantly different strategies compared to the conrol group. We evaluate the performance of AI agents under similar conditions.
Three scenarios are reported. In Scenario 14, participants experienced anger, and the five AI models consistently generate higher guesses (ranging from about 24.10 to 46.70) than they do in Scenario 16, the neutral control condition (where model means range from 7.20 to 37.02). These findings parallel the original conclusion that anger increases guesses (and thus reduces optimality). Once again, Llama performs differently with only a minor raise when playing against anger opponents. In contrast, Scenario 15 induces sadness; here, two of the models’ guesses (GPT-4o mini and Llama) are more close to the neutral condition, indicating no large deviations from control, consistent with Castagnetti et al. (2023) observation that sadness does not significantly alter players’ strategies. At the same time, Gemini Flash, Claude Sonnet, and GPT-4o react by increasing their numbers when playing against sad opponents. As in earlier replications, Llama’s guesses remain at the higher end in all emotional states, highlighting its tendency to overshoot relative to other large language models.
Scenario $n$ $p$ Function Opponents Paper result Model Model mean Model st. dev.
3 0.7 mean individuals experiencing anger Less optimal compared to control group Gemini Flash 24.10 15.51
Claude Sonnet 46.70 18.09
14 GPT-4o mini 33.20 7.89
GPT-4o 32.04 5.71
Llama 38.38 9.30
3 0.7 mean individuals experiencing sad emotions Sadness does not shift the players’ behavior Gemini Flash 25.88 15.18
Claude Sonnet 23.08 7.75
15 GPT-4o mini 21.18 9.30
GPT-4o 30.96 5.86
Llama 38.36 14.36
3 0.7 mean individuals experiencing neither anger nor sad emotions Lower numbers compared to players experiencing anger Gemini Flash 7.20 12.41
Claude Sonnet 18.76 8.67
16 GPT-4o mini 21.66 13.34
GPT-4o 22.2 11.84
Llama 37.02 10.78
Table 6: Replication results for Castagnetti et al. (2023) with a LLM player.
4 Discussion
In the Introduction, we formulated 7 questions that can help us to understand strategic reasoning of the artificial intelligence. Now we are ready to answer those questions based on the results of our Guess the number experiments.
Q1. Does LLM recognize the rules of the game and act in accordance with the rules? (Yes.)
In our dataset, 100% of LLM answers are legitimate. However, we had to modify our preliminary prompts several times by clarifying the expected format of the output to achieve this result. More vague initial versions of the prompt sometimes led to ambiguous answers such as intervals instead of a unique number. We think that a very detailed formulation of the task is a basic programming need and cannot be interpreted as a simplification of the problem for our AI agents.
Q2. Does LLM recognize the strategic context of the game? (Yes.)
We calculated the frequency of some keywords appearance in the text explanations of LLM players. More precisely, we look for the words ‘strategy’, ‘opponent’, ‘best response’, ‘dominant’, ‘Nash equilibrium’ that potentially indicate strategic nature of their considerations (see Table 7).
Model Strategy Opponent Best response Dominant Nash equilibrium
Gemini Flash 0.804 0.288 0.004 0.035 0.325
Claude Sonnet 0.474 0.074 0 0.001 0.146
GPT-4o mini 0.426 0.446 0.006 0.008 0.089
GPT-4o 0.442 0.384 0.026 0.008 0.238
Llama 0.576 0.335 0 0 0.021
Table 7: Frequency of keywords appearance in the text explanations of LLM players. For each LLM, the frequencies are calculated for all 16 scenarios together.
It appears that all LLMs actively use the word ‘strategy’ in their responses (the frequency varies between 42% and 80% depending on the model) and all LLMs except for Claude Sonnet use the word ‘opponent’ rather often (the frequency varies between 29% and 45%). Such game-theoretic concepts and notions as dominant strategies and best responses were not actively used by the models, while Gemini Flash, GPT-4o and Claude Sonnet mentioned Nash equilibrium from time to time. Note that the numbers in Table 7 can be considered as lower bounds for the share of strategic explanations. In practice, the models may use synonyms to express strategic ideas. For example, the following Claude Sonnet’s explanation does not mention the words ‘opponent’ or ‘best response’ but definitely provides an example of strategy optimization based on beliefs about the opponents’ strategies:
As someone with a low cognitive reflection test score, I tend to go with my first instinct without thinking too deeply about strategy. The game says to choose a number closest to 2/3 of the average, so I’ll pick a number that feels close to that. Most people will probably choose numbers around the middle, like 50. Two-thirds of 50 is about 33, so I’ll go with a number near that. I think 35 sounds good - it’s close to 33 but a little higher to be safe. I don’t want to overthink it, so I’ll stick with 35 as my final answer.
Frequent use of some of the keywords leads us to the conclusion that LLMs indeed apply strategic reasoning to derive the answer. The following explanation of Claude Sonnet reinforces our impression about the LLMs abilities to think strategically:
As a player with a high cognitive reflection test (CRT) score, I will approach this problem analytically. First, let’s consider the game theory aspects: 1. If all players chose randomly, the average would be 50, and 2/3 of that would be 33.33. 2. However, knowing this, rational players would choose lower numbers. 3. Anticipating this, even more strategic players would go even lower. Given that all players have high CRT scores, we can assume they will all follow this logic to multiple levels. This is known as the ’beauty contest’ game in game theory. In practice, studies have shown that even among highly analytical players, choices tend to converge around the second or third level of reasoning: - Level 0: Random choice (50) - Level 1: 2/3 of 50 $\approx$ 33 - Level 2: 2/3 of 33 $\approx$ 22 - Level 3: 2/3 of 22 $\approx$ 15 Considering that all players have high CRT scores, most will likely reach at least the second or third level of reasoning. However, to account for potential variations and to slightly undercut the most common strategic choice, I will choose 14. This number is slightly below the third level of reasoning (15), which gives me an edge if most players stop at that level. It’s also not so low that it would be far off if some players don’t reason as deeply or if there’s some distribution around the level 3 choice. Therefore, my chosen number is 14.
Q3. Are LLM’s decisions in line with the expected comparative statics with respect to the parameters of the experiment? (Yes, with rare exceptions mostly produced by Llama.)
All models considered in the paper demonstrate expected adaptability to the changing set of competitors in almost all experiments. Table 3 shows that the LLMs play higher numbers when maximum is the aggregate function. Four out of five LLMs play on average higher numbers when median is the aggregate function compared to the mean. The only exception is Llama that demonstrates the opposite trend. It follows from Table 4 that all models except Llama surely respect the sophistication level of the opponents by playing lower numbers against the participants of academic conferences than against the first year students. Surprisingly, in two-player experiments on average Llama play very similar numbers against students and professionals. With the exception of GPT-4o mini, all models play lower numbers against large groups of competing students than against a single opposing student which is in line with the economic intuition. Table 5 reveals that all models except Llama play lower numbers against players with high CRT score than against those with low CRT score. Once again, Llama behaves vice-versa in a game with $p=\frac{1}{2}$. Finally, Table 6 confirms our expectations that LLMs should play higher numbers when they play against individuals experiencing anger.
Q4. Can LLM find an analytical solution to the game? (Indefinite).
Based on the results of our experiments, we fail to give a definite answer to this question. On the one hand, for some models each appearance of the words ‘Nash equilibrium’ is accompanied by numerically correct strategy of 0. However, it is also often accompanied by considerations that are different from the definition of a Nash equilibrium. For example, consider the following response of GPT-4o in scenario 6:
In this game, each player needs to select a number between 0 and 100, with the goal of choosing a number that is closest to 2/3 of the average of both chosen numbers. In a purely rational setting with players having complete knowledge of game theory, the Nash equilibrium would drive the numbers towards 0. This occurs through iterative elimination of dominated strategies: if everyone initially chooses numbers randomly, the average would be 50, leading players to aim for 2/3 of 50, which is approximately 33.33. In subsequent iterations, players choose 2/3 of the previous target, continuing this process, which would eventually converge towards 0. However, our situation involves first-year undergraduate students without formal game theory training. These players are likely to focus on intuitive reasoning rather than fully iterated strategies. They might assume others will choose numbers that are higher than 0 and use simpler estimations than iterated eliminations to settle around lower numbers. Considering this context, a reasonable assumption is that these students will choose numbers between 20 and 30, as many will likely estimate 2/3 of a more central average around these numbers due to initial guesses hovering around the midpoint. Therefore, aiming lower than the naive initial target, but not as low as the ultimate Nash strategy, is likely advantageous. Choosing number 22 takes into account both rudimentary player expectations and the initial strategic logic of such a game. This choice balances between realistic player behavior and effective strategic positioning. Thus, my balanced choice in understanding how a non-game-theory-informed group might think is to select the number 22.
In this example, GPT-4o incorrectly applies the concept of a dominated strategy. Even more obvious mistake was made by Llama in the following response to scenario 7:
In this game, the goal is to choose a number that is closest to 2/3 of the average of all chosen numbers. To maximize the chances of winning, we should consider the possible strategies of the audience. Since we are in the same group as the audience, we can assume they will use the same strategy. One possible strategy is to use the ’Nash equilibrium’ approach, where each player chooses a number that is the expected value of the average, given the other players’ choices. However, this approach assumes that the other players are random and independent, which is not the case here. We are competing against a group of economics and psychology experts, who are likely to be aware of this strategy and may adapt to it. A more effective approach is to consider the ’median’ strategy, where each player chooses a number that is the median of the possible range. This strategy is more robust to the audience’s potential adaptations. Since we are choosing between 0 and 100, the median of the range is 50. However, to make our choice more effective, we should consider the fact that 2/3 of the average will likely be closer to the median than the extremes. Therefore, we should choose a number that is slightly above the median, but not too high. A good choice would be 50, but to be more strategic, we can choose a number that is slightly above 50, such as 50.5. However, since the audience may also consider this strategy, we should choose a number that is even more strategic, which is the number 50.
At the same time, there were also responses with correct identification and interpretation of Nash equilibrium. Consider the following Gemini’s answer in scenario 6:
This game is a fascinating example of how common sense reasoning can lead to a surprising outcome. Here’s how we can think about it: **The Intuitive Trap:** The natural inclination is to choose a number around 66, as 2/3 of 100 is roughly 66. However, this logic falls apart quickly. If everyone chooses 66, the average would be 66, and 2/3 of that is still 66, leading to a tie. **The Iterative Reasoning:** The key is to think about what your opponents are likely to do. They are likely to reason in the same way you are, realizing the trap of 66. So they might think to choose a lower number, say 50, hoping to shift the average down. But you’ll also realize this, and choose even lower, say 40, hoping to further shift the average. This iterative reasoning, where each player anticipates the other’s actions, will likely drive the chosen numbers lower and lower. **The Nash Equilibrium:** The logical end of this iterative process is that everyone chooses 0. In this scenario, the average is 0, and 2/3 of the average is still 0. While this might seem counterintuitive, it’s the only strategy that cannot be improved upon by any individual player, given what the others are doing. This is known as a Nash Equilibrium in game theory. **Therefore, I will choose the number 0.**
Q5. Can LLM correctly use the analytical solution when choosing their strategy? (No.)
Even if we do believe that LLMs know the concept of dominant strategies, they miserably fail to apply it properly. We learn from Table 4 that for the case $n=2$, average LLMs’ strategies are far from the dominant strategy of 00. We do not analyze the ability to play Nash equilibrium for $n>2$ because playing 0 probably leads to a defeat to real-world players who rarely play 0.
Q6. Do LLMs perform differently? (Yes.)
The most convincing argument against the idea that all LLMs produce almost the same outcome is the behavior of Llama. Llama systematically play much higher numbers than other LLMs (see Tables 26). Moreover, Llama often does not follow the expected comparative statics (see Q3), in contrast to other LLMs.
Q7. Are LLM’s strategies similar to strategies played by human players? (No in absolute values. Yes in strategic terms).
If one would cast a look at Tables 26, one could notice that the difference between the strategies played by humans and by LLMs is rather large. LLMs tend to overestimate the level of sophistication of their opponents and play lower numbers. The only exception is Llama that systematically plays high numbers that are often higher than the numbers played by humans. However, our results suggest that LLMs (except Llama) are rather successful in identifying the underlying strategic environment and adapting the strategies to the changing set of parameters of the game in the same way as human players do (see Q2 and Q3).
5 Conclusion
This study explored the performance of various LLMs in replicating human-like strategic reasoning across a range of experimental settings derived from five behavioral studies dealing with the Guess the number game. By replicating the experiments with such models as Claude Sonnet, Gemini Flash, GPT-4o, GPT-4o Mini, and Llama, we assessed their ability to align with human decision-making patterns, adapt to varying experimental parameters, and approximate theoretical predictions.
Our results show that 4 out of 5 modern LLMs in consideration (Claude Sonnet, Gemini Flash, GPT-4o, and GPT-4o Mini) systematically behave in a more sophisticated way compared to the participants of the original experiments. The LLMs play lower numbers compared to the numbers played by the human players in a similar setting. The only exception is Llama, it typically plays higher or nearly equal numbers compared to human players’ strategies. Also, Llama exhibited more static strategies, reflecting lower sensitivity to changes in experimental settings. All models failed to play a dominant strategy in a two-player game. These results suggest that while some LLMs are capable of emulating nuanced strategic behavior, their responses are often shaped by their underlying architectures and design priorities.
This study contributes to the growing body of research on the potential of LLMs to model human behavior in economic decision-making contexts. Future work could extend these analyses to more complex strategic environments, incorporate additional behavioral datasets, and explore ways to enhance the adaptability of LLMs to further bridge the gap between artificial and human decision-making. The convergence of AI’s calculated rationality with human intuition and behavior would open new avenues for enhancing predictive models and designing economic policies that account for the bounded rationality in human economic activities.
Acknowledgements
We would like to thank the participants of the European Meeting on Game Theory 2024 (SING19) and participants of the HSE International Laboratory of Game Theory and Decision Making research seminar for their valuable comments and suggestions.
Funding
The authors gratefully acknowledge support from the Basic Research Program of the HSE University. The funding source had no involvement in the research.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used GPT-4o, GPT-4o mini, Gemini-1.5-flash, Claude-3.5-Sonnet, and Llama-3.1-8B-Instruct-Turbo as economic agents in order to run the experiments. Also, GPT-4o was used for language proofreading purposes. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Data availability
The data supporting the findings of this study is publicly available in Harvard Dataverse: https://doi.org/10.7910/DVN/LURKWJ (accessible upon publication of the manuscript).

### v2

@@
 Abstract
-A Keynesian beauty contest is a wide class of games of guessing the most popular strategy among other players.
+A $p$-beauty contest is a wide class of games of guessing the most popular strategy among other players.
 In particular, guessing a fraction of a mean of numbers chosen by all players is a classic behavioral experiment designed to test iterative reasoning patterns among various groups of people.
@@
 Our results contribute to the discussion on the accuracy of modeling human economic agents by artificial intelligence.
-Keywords: beauty contest, guess the number, dominant strategy, AI, LLM
+Keywords: $p$-beauty contest, guess the number, dominant strategy, AI, LLM
 JEL codes: C99, D90, C72
@@
 Still, at this point, it is not clear to what extent LLM can simulate the human’s behavior.
-Several papers investigated the differences between LLM decisions and the participants of economic experiments.
-In the recent manuscript, Horton (2023) replicates several classic experiments with LLM players and advocates the use of LLMs as models for ordinary economic agents.
-Our paper contributes to this strand of literature by studying the behavior of LLMs in the classic Guess the number game.
-This game belongs to a wider class of Keynesian beauty contest experiments and is of particular importance because it is designed to study level-$k$ reasoning, the ability of a player to make a sequence of conclusions as a function of their level of sophistication.
-To what extent LLM behaves like a human in making sequential conclusions?
-To answer this question, we replicate a series of well-known experiments with human participants by asking LLMs to play against the same groups of competitors.
+The revelation of differences between the decisions of the participants of economic experiments and LLMs’ decisions in similar circumstances becomes an important challenge.
+Horton (2023) replicates several classic experiments with LLM players and advocates the use of LLMs as models for ordinary economic agents.
+Akata et al.
+(2023) reveal that LLMs underperform in the iterated games that require coordination such as the iterated Battle of Sexes.
+Brookins & DeBacker (2023) show that LLMs prefer fair decisions in a Dictator game, even more than human players do.
+Goli & Singh (2024) study the intertemporal preferences of LLMs and figure out that artificial players are less patient than humans.
+In general, different LLMs have different peculiarities: some are very sensitive to game structure, others have issues with context framing (Lorè & Heydari, 2024).
+Our paper contributes to this strand of literature by studying the behavior of LLMs in the classic Guess the number game which belongs to a wider class of $p$-beauty contest games.
+These $p$-beauty contest games are of particular importance because they arise in various industries where the profit of a firm or the payoff of an individual player depend on the median preferences of all economic agents.
+The behavior of short-term traders and the resulting asset pricing is regarded as a variant of a beauty contest game (Allen et al., 2006, Cespa & Vives, 2015).
+On the prediction markets, the desire to guess the most popular opinion is often considered as a behavioral driving force in addition to guessing the fundamental probabilities of an outcome (Marinovic et al., 2011).
+The beauty contest auction is an important alternative to price-only mechanisms of allocating goods to the bidders (Yoganarasimhan, 2016).
+Whereas algorithmic trading and betting strategies written by humans have long been used on the markets, the appearance of strategies generated by LLMs is a rather recent phenomenon.
+Investigation of the LLMs’ performance in a $p$-beauty contest game (both absolute and relative to humans) would help to predict their performance in the above-mentioned markets.
+The Guess the Number game tests the ability of a player to make a sequence of conclusions, and the outcome depends on their level of sophistication.
+The literature provides several important observations about the performance of heterogeneous players in various strategic settings.
+First, and most obvious, more advanced players reach better outcomes.
+The latter observation is especially notable in antagonistic pure games of skill.
+Charness (1981) revealed that the objective quality of moves selected by chess players increases with the player’s skill.
+Levitt et al.
+(2011) considered a sequential game Race to 100 which is a pure game of skill.
+They conclude that the ability to perform backward induction leads to better results in the Race to 100 game.
+Second (and this is less obvious), the players pay attention to the opponents’ quality.
+In the Eichberger et al.
+(2008) experimental study, most of the participants (72%) who play the simultaneous games of strategic complements or strategic substitutes feel that they can predict the actions of a game theorist better than the actions of a granny.
+The same large majority (72%) prefer to play against a game theorist rather than against a granny.
+In the Prisoner’s Dilemma game, the high-ability players who learn that their partner is a low-ability one, decrease the level of cooperation (Lambrecht et al., 2024).
+In antagonistic games, more advanced human players demonstrate greater adaptability to competing environment.
+For example, in the centipede game, stronger chess players exploit the non-equilibrium play of weaker players (Palacios-Huerta & Volij, 2009).
+However, Levitt et al.
+(2011) show that the ability to perform backward induction is not related to ability to stop early in the centipede game and assert that, instead of a centipede game, the pure games of skill should be used to test the role of the level of sophistication.
+We conclude that in many strategic environments, the skill of the players affects the outcome by allowing them to choose better strategies and/or to show greater adaptability to the opponents.
+To what extent do LLMs behave like a human?
+To answer this question, we replicate a series of well-known experiments with human participants playing the Guess the Number game by asking LLMs to play against the same groups of competitors.
 The rules of the Guess the number game are as follows.
@@
 In case of a tie, all tied players get the corresponding share of the prize.
-For $p\in(0,1)$, there is a unique Nash equilibrium in the model where all players choose 00.
+When all real numbers from $[0,100]$ are allowed, for $p\in(0,1)$ there is a unique Nash equilibrium in the model where all players choose 00.
 In a particular case of $n=2$, choosing 00 is a weakly dominant strategy.
+If only integer numbers are allowed, there do exist additional Nash equilibria where the players play higher numbers.
+For example, if $p=0.5$, all strategy profiles where most of the players play $1$ and other players play 00, are additional Nash equilibria.
+If $p=2/3$, the profile $(1,\ldots,1)$ is the only additional Nash equilibrium.
 Multiple experiments show that in the Guess the number game people, in general, do not play Nash equilibrium.
@@
 However, in the case of professionals, this difference is not statistically significant (Grosskopf & Nagel, 2008).
+Rydval et al.
+(2009) strengthen this finding by identifying that only nearly 1/3 of all participants think in terms of dominant strategies in 2- or 3-person Guess the number-like games with dominant strategies, while 2/3 of all participants fail to identify the strategic properties of the game.
+In our study, we ask whether LLMs are able to identify the strategic nature of the game.
 Several theoretical models explaining the non-equilibrium behavior were proposed in the literature.
@@
 Cognitive ability is also an important determinant of the outcome of the Guess the number game.
-Brocas & Carrillo (2020) designed a variant of the Guess the number game and demonstrated that the equilibrium behavior increases significantly between 5 and 10 years of age (from 17.9% to 61.4%) and stabilizes afterwards.
-Back to the claasic variant of the game, players with higher scores in a cognitive ability test choose lower numbers (Burnham et al., 2009).
+Higher cognitive ability may manifest itself through better inductive reasoning, iterative dominance, and level-$k$ thinking (Carpenter et al., 2013).
+Brocas & Carrillo (2020) designed a variant of the Guess the number game and demonstrated that the equilibrium behavior increases significantly between 5 and 10 years of age (from 17.9% to 61.4%) and stabilizes afterward.
+Back to the classic variant of the game, players with higher scores in a cognitive ability test choose lower numbers (Burnham et al., 2009) and show faster convergence to equilibrium in iterated experiments (Gill & Prowse, 2016).
 Mixed evidence was reported in Brañas-Garza et al.
 (2012): the better performance in the CRT test that measures cognitive reflectiveness is associated with lower numbers in the Guess the number game, whereas the outcome of the Raven test measuring visual reasoning and analytic intelligence surprisingly was not associated with the successful performance in the Guess the number game.
+Interestingly, high cognitive ability test scorers better respond to the cognitive ability of their opponents (Gill & Prowse, 2016), whereas players whose abilities are below a certain threshold do not adapt strategically to the opponents’ level of sophistication at all Fehr & Huck (2016).
+The recent Gill et al.
+(2025) study distinguishes between cognitive ability and judgment.
+The authors show that whereas high cognitive ability shifts the strategy towards 0, high judgment subjects are less inclined to choose 0, even though their choices on average are lower than low judgment subjects.
 Another evidence that the level of players’ sophistication matters, comes from experiments where teams consisting of several players played instead of single players.
@@
 At the same time, teams consisting of 3 and 4 players outperform individual players (Kocher & Sutter, 2005, Sutter, 2005).
+One should also distinguish between cognitive abilities and cognitive effort.
+Alós-Ferrer & Buckenmaier (2021) propose an experiment where the deliberation time in the Guess the number game serves as a proxy for cognitive effort.
+The authors find evidence that longer deliberation time is related to playing strategies associated with higher numbers of reasoning steps.
+After running experiments with two families of games, Georganas et al.
+(2015) warn us that strategic sophistication varies across different families of games which prevents us from making too strong conclusions and expanding them from the Guess the number game to more general environments.
 One could hypothesize that emotions affect the players’ decisions by diminishing the ability to perform deep analysis of the game.
@@
 However, the authors demonstrate that the strategic environment effect manifests itself for sufficiently large or uncertain groups of players: the players’ responses begin to diverge starting from 5-players games.
+There are two principal dimensions of the problem in consideration.
+  1. 1.
+    For a wide range of parameters of a one-shot game, one can compare the strategies of humans and LLMs in a one-shot Guess the number game.
+  2. 2.
+    For a specific set of parameters, one can compare the strategies of humans and LLMs in the iterated version of the Guess the number game.
+While the first approach focuses on the role of changing conditions of the game, the second approach studies learning effects.
+Our paper follows the former one.
+We do not observe learning effects neither in humans nor in LLMs.
+In order to make a fair comparison, we took only the first iterations of the game from the previous literature.
+Similarly, we never asked LLMs to play the Guess the number game several times in a row during one iteration of an experiment, and we never provided any feedback to LLMs.
+In contrast, the working paper Lu (2024) follows the second approach.
 For more experimental and theoretical results on Guess the number games, we refer the reader to one of the surveys (Nagel, 2008, Nagel et al., 2017).
-Testing LLMs with the Guess the number game offers a dual benefit.
-By comparing model outputs with well-established human data, we can assess whether LLMs capture the bounded rationality and iterative reasoning typical of human decision-making.
-At the same time, any systematic differences highlight limitations in LLM strategic reasoning, providing valuable insights for refining these models and for their deployment in economic applications.
+Fan et al.
+(2024) presents a detailed overview of LLMs research published by 2023.
 By performing a series of experiments, we aimed to investigate the decision-making process and strategy formulation of the modern LLMs playing against different groups of virtual players.
@@
 Recall that some of the previous experiments included those participants who are definitely unfamiliar with dominant strategies and Nash equilibrium concepts, whereas other experiments included game theorists who are presumably able to find the theoretical equilibrium in the Guess the number game.
-Q5. Can LLM correctly use the analytical solution when choosing their strategy?
+Q5. Can LLM correctly use the analytical solution when choosing its strategy?
 In case of a positive answer, for the game of $n=2$ players we would expect that LLM always plays a weakly dominant strategy.
@@
 Q7. Are LLM’s strategies similar to strategies played by human players?
-Though this question is the most important for us, the nuances that may arise in questions Q1Q6 could affect the interpretation of the results.
+Though this question is the most important for us, the nuances that may arise in questions Q1–Q6 could affect the interpretation of the results.
 Therefore, we do not restrict our attention to Q7 solely.
+Testing LLMs with the Guess the number game offers a dual benefit.
+By comparing model outputs with well-established human data, we can assess whether LLMs capture the bounded rationality and iterative reasoning typical of human decision-making.
+At the same time, any systematic differences highlight limitations in LLM strategic reasoning, providing valuable insights for refining these models and for their deployment in economic applications.
 We experimented with the main LLMs that were available in 2024: GPT-4o, GPT-4o-Mini, Gemini-1.5-flash, Claude-3.5-Sonnet (20240620), Llama-3.1-8B-Instruct-Turbo.
@@
 For example, in a money request game LLMs are found to be less sophisticated than human players (Gao et al., 2024).
-After the initial publication of our working paper (Dagaev et al., 2024), we found several new relevant working papers that we would like to mention.
-Lu (2024) also runs beauty contest experiments with AI players.
-Whereas the latter paper focuses on the iterated standard beauty contest game and dynamics of the LLMs decisions, our paper replicates a wide class of one-shot experiments with different groups of competitors.
-Guo et al.
-(2024) introduced EconArena environment that allows to organize various competitions between AI models.
+Finally, note that we did not organize games between LLMs in this study.
+We refer those who are interested in tournaments between artificial players to Guo et al.
+(2024) who introduced EconArena environment that allows to organize various competitions between AI models.
 A beauty contest game is one of the first games that was implemented on EconArena.
@@
 Retrieved December 22, 2024.
+For reproducibility purposes we have now detailed all relevant experimental parameters in Table A1 in the Appendix.
+Specifically, we employed the default settings provided by the respective APIs for each model version—namely, GPT-4o, GPT-4o mini, Gemini-1.5-flash, Claude-3.5-Sonnet, and Llama-3.1-8B-Instruct-Turbo—including default token limits and temperature settings.
+The data collection process was repeated within a single session environment until 50 valid responses were obtained.
 The experiments were structured into 16 distinct scenarios, each characterized by a combination of factors including the type of aggregate statistic used to determine the winning number (Function), the target percentage ($p$) of the aggregate statistic, the number of players involved ($n$), and the composition of the opponent group (Opponents).
@@
 9 18 2/3 Mean Audience of game theory conferences
-10 Brañas-Garza et al. (2012) 24 1/2 Mean individuals with high CRT score
+10 Braùas-Garza et al. (2012) 24 1/2 Mean individuals with high CRT score
 11 24 1/2 Mean individuals with low CRT score
@@
 Table 4: Replication results for Grosskopf & Nagel (2008) with a LLM player.
-Replication results for Brañas-Garza et al.
+Replication results for Braùas-Garza et al.
 (2012) experiment are reported in Table 5.
@@
 Llama 41.06 12.30
-Table 5: Replication results for Brañas-Garza et al. (2012) with a LLM player.
+Table 5: Replication results for Braùas-Garza et al. (2012) with a LLM player.
 Table 6 replicates the experiment from Castagnetti et al.
@@
 **Therefore, I will choose the number 0.**
-Q5. Can LLM correctly use the analytical solution when choosing their strategy? (No.)
+Q5. Can LLM correctly use the analytical solution when choosing its strategy? (No.)
 Even if we do believe that LLMs know the concept of dominant strategies, they miserably fail to apply it properly.
@@
 The most convincing argument against the idea that all LLMs produce almost the same outcome is the behavior of Llama.
-Llama systematically play much higher numbers than other LLMs (see Tables 26).
+Llama systematically play much higher numbers than other LLMs (see Tables 2 –6).
 Moreover, Llama often does not follow the expected comparative statics (see Q3), in contrast to other LLMs.
 Q7. Are LLM’s strategies similar to strategies played by human players? (No in absolute values. Yes in strategic terms).
-If one would cast a look at Tables 26, one could notice that the difference between the strategies played by humans and by LLMs is rather large.
+If one would cast a look at Tables 2 –6, one could notice that the difference between the strategies played by humans and by LLMs is rather large.
 LLMs tend to overestimate the level of sophistication of their opponents and play lower numbers.
@@
 These results suggest that while some LLMs are capable of emulating nuanced strategic behavior, their responses are often shaped by their underlying architectures and design priorities.
+A limitation of our study is that we did not incorporate an explicit payoff structure in the AI prompts.
+The original experiments featured varying incentive mechanisms, which may have influenced participants’ choices.
+Our decision to leave incentives unspecified—while aiming to replicate original conditions—might have affected the observed strategic behavior of the LLMs.
+Future research should investigate how different incentive specifications influence AI decision-making in strategic settings.
+Another limitation is that the alternative to use the same instructions that humans got in the respective experiments, is not feasible.
+The context matters.
+For example, participants of a game theory conference understand who they are.
+In contrast, LLMs need additional information about their competitors and themselves in the prompt.
+Therefore, we stick to the second-best option of using prompts that are as similar to the original instructions as possible.
 This study contributes to the growing body of research on the potential of LLMs to model human behavior in economic decision-making contexts.
@@
 Data availability
-The data supporting the findings of this study is publicly available in Harvard Dataverse: https://doi.org/10.7910/DVN/LURKWJ (accessible upon publication of the manuscript).
+The data supporting the findings of this study is publicly available in Harvard Dataverse: https://doi.org/10.7910/DVN/LURKWJ.